OpenStudy (anonymous):

Find the volume of a sphere with a radius of 2.8 meters. Round to the nearest whole number. Use = 3.14.

OpenStudy (jlastino):

The Volume of a Spehere = (4/3) pi r^3 Sub r= 2.8 pi= 3.14

OpenStudy (jlastino):

*Sphere

jhonyy9 (jhonyy9):

will be 4/3 pi *21,952 =29,269 pi m^3